A nozzle is a device for increasing the velocity of a steadily flowing stream. At the inlet to




a certain nozzle, the enthalpy of the fluid passing is 3000 kJ/kg and the velocity is 60 m/s.




At the discharge end, the enthalpy is 2762 kJ/kg. The nozzle is horizontal and there is




negligible heat loss from it. (a) Find the velocity at exists from the nozzle. (b) If the inlet




area is 0.1 m2 and the specific volume at inlet is 0.187 m3




/kg, find the mass flow rate.




(c) If the specific volume at the nozzle exit is 0.498 m3




/kg, find the exit area of the nozzle.

1
Expert's answer
2021-11-15T17:55:09-0500

a)

"H_1v_1=H_2v_2,\implies v_2=v_1\frac{H_1}{H_2}=65~\frac ms,"

b)

"\dot m=\frac{S_1v_1}{\rho_1}=32~\frac{kg}s,"

c)

"\dot m=\frac{S_2v_2}{\rho_2},\implies S_2=\frac{\dot m\rho_2}{v_2}=1~m^2."